# Ordering kinetics in the active Ising model

### v1

@@ Abstract
We undertake a numerical study of the ordering kinetics in the two-dimensional ($2d$) active Ising model (AIM), a discrete flocking model with a non-conserved scalar order parameter. We find that for a quench into the liquid-gas coexistence region and in the ordered liquid region, the characteristic length scale of both the density and magnetization domains follows the Lifshitz-Cahn-Allen (LCA) growth law: $R(t)\sim t^{1/2}$, consistent with the growth law of passive systems with scalar order parameter and non-conserved dynamics. The system morphology is analyzed with the two-point correlation function and its Fourier transform, the structure factor, which conforms to the well-known Porod’s law, a manifestation of the coarsening of compact domains with smooth boundaries. We also find the domain growth exponent unaffected by different noise strengths and self-propulsion velocities of the active particles. However, transverse diffusion is found to play the most significant role in the growth kinetics of the AIM. We extract the same growth exponent by solving the hydrodynamic equations of the AIM.
@@ I Introduction
Vicsek and collaborators years ago introduced a minimal model Vicsek et al. (1995) of active particles that move with a constant speed and orient via a ferromagnetic interaction with a neighborhood similar to the XY model. In the Vicsek model (VM) activity can stabilize the ordered phase even in two dimensions which is not possible in the $2d$ XY model where the long-range fluctuation destroys the ordered phase following Mermin-Wagner theorem  Toner and Tu (1995, 1998). Then, about a decade ago, Solon and Tailleur introduced the active Ising model (AIM) Solon and Tailleur (2013, 2015) where the continuous rotational symmetry of the VM is replaced by a discrete symmetry. In the AIM, each particle assumes two possible states allowing the particle to propel in a preferred direction which changes upon interaction with other particles at the same lattice site. The AIM retains the essential part of the VM physics and exhibits flocking behavior with three different phases at steady state: disordered gas at high noise and low densities, polar liquid at low noise and high densities, and a phase-separated liquid-gas coexistence state at intermediate densities. However, a key difference between the VM and the AIM arises in the steady-state behavior of the coexistence region. In this region, AIM shows a macrophase separation associated with normal density fluctuations whereas the VM is characterized by a microphase separation with giant density fluctuations. The flocking transition in the AIM is a first-order liquid-gas phase transition similar to the VM, however, for zero activity, despite the dynamics being non-equilibrium, the AIM shows a second-order phase transition belonging to the Ising universality class.
@@
Although significant progress has been made to understand the steady state properties of various active systems Toner et al. (2005); Giomi et al. (2013); Cates and Tailleur (2015); Solon et al. (2015); Siebert et al. (2018); Wysocki and Rieger (2020); Mangeat et al. (2020); Chatterjee et al. (2020a); Kürsten and Ihle (2020); Fruchart et al. (2021); Solon et al. (2022); Chatterjee et al. (2022); Codina et al. (2022); Chatterjee et al. (2023); Karmakar et al. (2023), there is much to explore in the realm of ordering kinetics in active systems that relaxes to a non-equilibrium steady state (NESS). Understanding the intrinsic non-equilibrium dynamics that drive an active system towards its steady state is of fundamental as well as practical relevance. Unlike active systems, ordering kinetics in non-equilibrium passive systems have been studied over several decades Bray (2002, 1994); Puri (2009); Ahmad et al. (2012); Shrivastav et al. (2014); Kumar et al. (2017); Chatterjee et al. (2020b). Domain growth in passive systems with non-conserved scalar order parameters follows the Lifshitz-Cahn-Allen (LCA) growth law: $R(t)\sim t^{1/2}$ (Model A of order-parameter kinetics) whereas passive systems with conserved order parameter follow a Lifshitz-Slyozov-Wagner (LSW) growth law: $R(t)\sim t^{1/3}$ (Model B of order-parameter kinetics), $R$ being average size of domains. Using tools that quantify the kinetics of passive systems, several active systems have been explored. These include Active Model B Wittkowski et al. (2014); Pattanayak et al. (2021a, b), active nematics Mishra et al. (2014), self-propelled particles in disordered medium Das et al. (2018), Model B with nonreciprocal activity Saha et al. (2020), Kuramoto oscillators Rouzaire and Levis (2022), active Brownian particles Dittrich et al. (2023) and motility-induced phase separated (MIPS) clusters Caporusso et al. (2023). Moreover, an interesting observation of multiple coarsening length scales was made in the prototypical VM Katyal et al. (2020) where velocities are found to align over a faster-growing length scale compared to density. Another intriguing result of an active system with a non-conserved vector order parameter following the growth law of the non-conserved scalar order parameter field has also been observed Dikshit and Mishra (2023). Since AIM is a minimal flocking model with a rich phase behavior, studying the growth kinetics of this model will allow us to interpret the origin of large flocks in terms of microscopic interactions.
@@
In this paper, we explore the phase ordering kinetics of the AIM with a non-conserved scalar order parameter. Quenching the AIM inside the spinodal region results in the formation of small positively or negatively magnetized clusters which in the late stage of the coarsening merge to form a single, macroscopic domain of one spin type Solon and Tailleur (2015). A few questions arise in this context: (a) Does the domain morphology follow the same pattern and growth law for quenches into the coexistence and in the ordered liquid region? (b) Since the order parameter is non-conserved in AIM, how does the growth law relate to the established growth law of similar passive systems? (c) Do the density and magnetization align over the same length scale? Katyal et al. (2020) (d) What is the impact of noise and particle activity on the domain growth? and (e) What is the role of diffusion in the domain growth dynamics? We address these issues by analyzing the ordering dynamics of the $2d$ AIM on a square lattice via Monte Carlo simulations and by solving the AIM hydrodynamic equations using the finite difference method.
@@ II Model
We consider $N$ particles on a two-dimensional square lattice $L\times L$ with periodic boundary conditions. Thus the average particle density is $\rho_{0}=N/L^{2}$. Each lattice site $i$ can accommodate an arbitrary number of particles $n_{i}^{\sigma}$ with spin $\sigma=\pm 1$. Defining local density $\rho_{i}=n_{i}^{+}+n_{i}^{-}$ and magnetization $m_{i}=n_{i}^{+}-n_{i}^{-}$ we note that $\rho_{i}$ has no upper bound, while $m_{i}$ is bounded by $\rho_{i}$: $-\rho_{i}\leq m_{i}\leq\rho_{i}$. Each particle with a given spin state $\sigma$ can either flip to $-\sigma$ or jump to a nearest-neighbor site probabilistically.
The flipping rates are derived from a local ferromagnetic Potts Hamiltonian Mangeat et al. (2020); Chatterjee et al. (2020a) defined as:
$H_{i}=-\frac{J}{2\rho_{i}}\sum_{k=1}^{\rho_{i}}\sum_{l\neq k}\left(q\delta_{\sigma_{i}^{k},\sigma_{i}^{l}}-1\right)\,,$ (1)
where $J$ is the coupling between any two particles at site $i$ and the Kronecker delta $\delta_{\sigma_{i}^{k},\sigma_{i}^{l}}$ survives only for $\sigma_{i}^{k}=\sigma_{i}^{l}$. Eq. (1) with $q=2$ is the local Hamiltonian defined for the AIM Solon and Tailleur (2015). Local interaction implies that a particle can align with the average direction of all other particles at the same site.
The spin-flip transition rates are derived from the Potts Hamiltonian $H_{\rm APM}=\sum_{i}H_{i}$ according to the energy difference before and after the spin flip. For $q=2$, a particle with spin $\sigma$ flips its state according to the transition rate  (Mangeat et al., 2020; Chatterjee et al., 2020a):
@@
where $\gamma$ is the rate of particle flipping when $\sigma m_{i}=1$. In this paper, we choose $J=1$, and $\gamma=1$ without any loss of generality. The parameter $\beta$, denoted as “inverse temperature”, $\beta=T^{-1}$, in passive systems, controls the flip noise strength. Although the system under consideration is athermal, we denote the parameter $T$ as “temperature” from now on.
Moreover, each particle performs a biased diffusion on the lattice depending on the spin state $\sigma$. For $q=2$ particles perform a one-dimensional biased hopping along the ±$\pm x$-direction with the following hopping rate Mangeat et al. (2020); Chatterjee et al. (2020a):
$W_{\rm hop}(\sigma,p)=D\left[1+\epsilon({2\delta_{\sigma,p}-1})\right]\,,$ (3)
where $p$ denotes the direction of bias-hopping ($p=+1$ along the $+x$ direction and $p=-1$ along the $-x$ direction). The presence of other particles does not influence hopping rates and hence is independent of particle density. The hopping rate $D=1$ is constant along the upward and downward directions (±$\pm y$-direction). The parameter $\epsilon\in[0,1]$ $(0\leq\epsilon\leq q-1)$ controls the asymmetry between the purely diffusive limit $\epsilon=0$ and the purely ballistic limit $\epsilon=1$, while $D$ controls the overall hopping rate. On average, a particle drifts with speed $v=2D\epsilon$ in the direction set by the sign of its spin state (where lattice spacing is 1), while the total hopping rate $4D$, remains constant.
@@ III Simulation details
where $f(x)$ is a time-independent scaling function. $R(t)$, estimated from the decay of $C(r,t)$ generally show a power-law growth Puri (2009); Bray (2002, 1994):
@@
For scalar order parameters like the density field, the short-distance (large-$k$) behavior of the structure factor scaling function is given by the Porod’s law (for domains with smooth boundaries or scattering off sharp domain interfaces) which corresponds to $g(k)\sim k^{-(d+1)}$. Next, we present results for model parameters set by the average particle density $\rho_{0}$, temperature $T=\beta^{-1}$, self-propulsion velocity $\epsilon$, and diffusion constant $D$.
@@ IV.1 Phase diagram and domain morphology
The steady-state behavior of the AIM is summarized in the temperature-density ($T-\rho_{0}$) [Fig. 1(a)] and velocity-density ($\epsilon-\rho_{0}$) [Fig. 1(b)] phase diagrams. The general structure of the phase diagrams consists of a gas phase (G), a liquid phase (L), and a liquid-gas coexistence region (G+L) separated by the gas and liquid binodals. Qualitatively similar results were obtained earlier in Ref. Solon and Tailleur (2015). However, we obtain a critical temperature $T_{c}\simeq 2$ (above which no phase separation occurs irrespective of the density, critical density $\rho_{c}=\infty$) twice as large as in Ref. Solon and Tailleur (2015). This can be understood if one compares the flipping rate of Ref. Solon and Tailleur (2015), $W_{\rm flip}(\sigma\rightarrow-\sigma)=\gamma\exp\left(-\sigma\beta\frac{m_{i}}{\rho_{i}}\right)$ with Eq. (2) for $J=1$ which denotes that the effective $\beta$ considered in this paper is roughly twice as large as the $\beta$ of Ref. Solon and Tailleur (2015). In the $\epsilon-\rho_{0}$ phase diagram, the two binodals converge at a critical density $(\rho_{c}\simeq 2.9)$ for vanishing self-propulsion ($\epsilon=0$) which signifies a second-order phase transition belonging to the Ising universality class Solon and Tailleur (2013).
@@
To quantify the coarsening dynamics we performed simulations on a square lattice of size $400^{2}$ with periodic boundary conditions applied on both sides. Following quench at time $t=0$, the system is evolved up to $t=10^{5}$ using the MC algorithm described in Sec. III. All numerical data presented here are averaged over at least 300 independent realizations.
Figure 1: (Color online) Phase diagrams of the AIM showing the quench directions at fixed densities. (a) $(T,\rho_{0})$ phase diagram for self-propulsion velocity $\epsilon=1$. (b) $(\epsilon,\rho_{0})$ phase diagram at fixed temperature $T=0.9$. G and L denote the disordered gas and polar liquid regions whereas G+L denotes the phase-coexistence region. The blue stars indicate the initial high-temperature points in the phase diagram from where the system is quenched (indicated by arrows) into the coexistence region ($\rho_{0}=3$) and to the liquid region ($\rho_{0}=10$) [red stars]. $\rho_{\rm gas}$ and $\rho_{\rm liquid}$ are the gas and liquid binodal, respectively.
@@
A typical simulation of coarsening dynamics starts with a homogeneous initial configuration where particles with $\sigma=\pm 1$ are distributed randomly on each lattice site. This can also be described as an equilibrium configuration at infinite temperature because all configurations are equally likely. Then, after a quench well below the critical temperature ($T_{c}\simeq 2$), the homogeneous initial configuration starts evolving in time, and subsequent dynamics are governed by the formation and growth of $\sigma=1$ and $\sigma=-1$ rich domains. Such a time evolution of the local magnetization field is shown in Fig. 2 for $\epsilon=0$ and $\epsilon=1$. In the latter scenario, time evolution has been shown for a quench in both the coexistence (middle row) and liquid regions (bottom row). The non-equilibrium steady states (NESS) are characterized by a single band (at lower density) and a polar liquid (at higher density).
@@
The top panel, Fig. 2(a–e), depicts the ordering kinetics of a purely diffusive $(\epsilon=0)$ polar liquid at $\rho_{0}=5$. The domain morphology with increasing time exhibits a close resemblance with the passive Ising model Majumder and Das (2011). In the Ising model, the driving force for the domain growth is the curvature of the domain wall since the system surface energy can only decrease through a reduction in the total net surface area. In the $\epsilon=0$ limit, particles do not form high-density domains due to the diffusive movement of particles. Therefore, density-wise, the system remains homogeneous as we observe a steady growth of small, high-curvature to large, low-curvature domains. This is unsurprising as the $2d$ AIM for $\epsilon=0$ belongs to the same universality class of the passive $2d$ Ising model Solon and Tailleur (2015).
@@
Next, we looked at the evolution of AIM with self-propulsion velocity ($\epsilon=1$) quenching the system inside the spinodal [Fig. 2(f–j)] and homogeneous ordered regions [Fig. 2(k–o)]. The average densities representing these two regions correspond to $\rho_{0}=3$ and 10, respectively. Inside the spinodal region, the growth dynamics are driven by spinodal decomposition and result in the formation of numerous small clusters of negative and positive spins $(t=10^{2})$. The coarsening then stems from the merging of these clusters $(t=10^{3}$ and $t=10^{4})$, until a single, macroscopic domain emerges $(t=10^{5})$ in the steady state. A quench deep inside the homogeneous ordered region also results in similar dynamics of cluster formation and coalescence of those clusters into a single large liquid domain (in this paper, for $\epsilon>0$, the word cluster is used interchangeably with domain). With extreme self-propelled particles ($\epsilon=1$), although the high-density clusters are strongly biased along the horizontal directions, they can also grow along the transverse direction due to the constant transverse diffusion $D$. When these clusters merge into a larger cluster, it always tries to minimize the surface energy by decreasing the surface area. Accordingly, the coarsening process leads to a single band [Fig. 2(j)] with domain walls having the lowest curvature (the curvature of a straight line is zero but a liquid band with a perfectly straight domain boundary can only happen when there is no thermal fluctuation). The focus of this study is, therefore, to analyze the coarsening dynamics shown in Fig. 2 which we will do next.

### v2

@@ Abstract
We undertake a numerical study of the ordering kinetics in the two-dimensional ($2d$) active Ising model (AIM), a discrete flocking model with a conserved density field coupled to a non-conserved magnetization field. We find that for a quench into the liquid-gas coexistence region and in the ordered liquid region, the characteristic length scale of both the density and magnetization domains follows the Lifshitz-Cahn-Allen (LCA) growth law: $R(t)\sim t^{1/2}$, consistent with the growth law of passive systems with scalar order parameter and non-conserved dynamics. The system morphology is analyzed with the two-point correlation function and its Fourier transform, the structure factor, which conforms to the well-known Porod’s law, a manifestation of the coarsening of compact domains with smooth boundaries. We also find the domain growth exponent unaffected by different noise strengths and self-propulsion velocities of the active particles. However, transverse diffusion is found to play the most significant role in the growth kinetics of the AIM. We extract the same growth exponent by solving the hydrodynamic equations of the AIM.
@@ I Introduction
Vicsek and collaborators years ago introduced a minimal model Vicsek et al. (1995) of active particles that move with a constant speed and orient via a ferromagnetic interaction with a neighborhood similar to the XY model. In the Vicsek model (VM) activity can stabilize the ordered phase even in two dimensions which is not possible in the $2d$ XY model where the long-range fluctuation destroys the ordered phase following Mermin-Wagner theorem  Toner and Tu (1995, 1998). Then, about a decade ago, Solon and Tailleur introduced the active Ising model (AIM) Solon and Tailleur (2013, 2015) where a discrete symmetry replaces the continuous rotational symmetry of the VM. In the AIM, each particle assumes two possible states allowing the particle to propel in a preferred direction which changes upon interaction with other particles at the same lattice site. The AIM retains the essential part of the VM physics and exhibits flocking behavior with three different phases at steady state: disordered gas at high noise and low densities, polar liquid at low noise and high densities, and a phase-separated liquid-gas coexistence state at intermediate densities. However, a key difference between the VM and the AIM arises in the steady-state behavior of the coexistence region. In this region, AIM shows a macrophase separation associated with normal density fluctuations whereas the VM is characterized by a microphase separation with giant density fluctuations. The flocking transition in the AIM is a first-order liquid-gas phase transition similar to the VM, however, for zero activity, despite the dynamics being non-equilibrium, the AIM shows a second-order phase transition belonging to the Ising universality class.
@@
Although significant progress has been made to understand the steady state properties of various active systems Toner et al. (2005); Giomi et al. (2013); Cates and Tailleur (2015); Solon et al. (2015a); Siebert et al. (2018); Wysocki and Rieger (2020); Mangeat et al. (2020); Chatterjee et al. (2020a); Kürsten and Ihle (2020); Fruchart et al. (2021); Solon et al. (2022); Chatterjee et al. (2022); Codina et al. (2022); Chatterjee et al. (2023); Karmakar et al. (2023), there is much to explore in the realm of ordering kinetics in active systems that relaxes to a non-equilibrium steady state (NESS). Understanding the intrinsic non-equilibrium dynamics that drive an active system towards its steady state is of fundamental as well as practical relevance. Unlike active systems, ordering kinetics in non-equilibrium passive systems have been studied over several decades Bray (2002, 1994); Puri (2009); Ahmad et al. (2012); Shrivastav et al. (2014); Kumar et al. (2017); Chatterjee et al. (2020b). Domain growth in passive systems with non-conserved scalar order parameters follows the Lifshitz-Cahn-Allen (LCA) growth law: $R(t)\sim t^{1/2}$ (Model A of order-parameter kinetics) whereas passive systems with conserved order parameter follow a Lifshitz-Slyozov-Wagner (LSW) growth law: $R(t)\sim t^{1/3}$ (Model B of order-parameter kinetics), $R(t)$ being average size of domains. However, domain growth in Model C systems where a non-conserved scalar order parameter is coupled to a conserved density field (similar to the AIM), is governed by the growth exponent of either Model A or Model B where it depends on the quenching regime. Quench into the order-disorder coexistence region leads to a $R(t)\sim t^{1/3}$ growth whereas the asymptotic growth law for a quench into the ordered region is $R(t)\sim t^{1/2}$ Kockelkoren and Chaté (2002).
Utilizing tools that quantify the ordering kinetics of passive systems, several active systems have been explored. These include Active Model B Wittkowski et al. (2014); Pattanayak et al. (2021a, b), active nematics Mishra et al. (2014), self-propelled particles in disordered medium Das et al. (2018), Model B with nonreciprocal activity Saha et al. (2020), Kuramoto oscillators Rouzaire and Levis (2022), active Brownian particles Dittrich et al. (2023) and motility-induced phase separated (MIPS) clusters Caporusso et al. (2023). Moreover, an interesting observation of multiple coarsening length scales was made in the prototypical VM Katyal et al. (2020) where velocities are found to align over a faster-growing length scale compared to density. Another intriguing result of an active system with a non-conserved vector order parameter following the growth law of the non-conserved scalar order parameter field has also been observed Dikshit and Mishra (2023). Since AIM is a minimal flocking model with a rich phase behavior, studying the growth kinetics of this model will allow us to interpret the origin of large flocks in terms of microscopic interactions.
@@
In this paper, we explore the phase ordering kinetics of the AIM. Quenching the AIM inside the spinodal region results in the formation of small positively or negatively magnetized clusters which in the late stage of the coarsening merge to form a single, macroscopic domain of one spin type Solon and Tailleur (2015). A few questions arise in this context: (a) Does the domain morphology follow the same pattern and growth law for quenches into the coexistence and in the ordered liquid region? (b) Since the AIM possesses both conserved and non-conserved order parameters, how does the AIM growth law relate to the established growth law of similar passive systems? (c) Do the density and magnetization correlate over the same length scale? Katyal et al. (2020) (d) What is the impact of noise and particle activity on the domain growth? and (e) What is the role of diffusion in the domain growth dynamics? We address these issues by analyzing the ordering dynamics of the $2d$ AIM on a square lattice via Monte Carlo simulations and by solving the AIM hydrodynamic equations using the finite difference method.
@@ II Model
We consider $N$ particles on a two-dimensional square lattice $L\times L$ with periodic boundary conditions. Thus the average particle density is $\rho_{0}=N/L^{2}$. Each lattice site $i$ can accommodate an arbitrary number of particles $n_{i}^{\sigma}$ with spin $\sigma=\pm 1$. Defining local density $\rho_{i}=n_{i}^{+}+n_{i}^{-}$ and magnetization $m_{i}=n_{i}^{+}-n_{i}^{-}$, we note that $\rho_{i}$ has no upper bound, while $m_{i}$ is bounded by $\rho_{i}$: $-\rho_{i}\leq m_{i}\leq\rho_{i}$. Each particle with a given spin state $\sigma$ can either flip to $-\sigma$ or jump to a nearest-neighbor site probabilistically.
The flipping rates are derived from a local ferromagnetic Ising Hamiltonian defined as Solon and Tailleur (2013, 2015):
$H_{i}=-\frac{J}{2\rho_{i}}\sum_{j}\sum_{k\neq j}\sigma_{i}^{j}\sigma_{i}^{k}\,,$ (1)
where $J$ is the coupling between any two particles at site $i$. Local interaction implies that a particle can align with the average direction of all other particles at the same site. The Hamiltonian of Eq. (1) can be rewritten as:
$H_{i}=-\frac{J}{2}\left(\frac{m_{i}^{2}}{\rho_{i}}-1\right)\,.$ (2)
When a particle in spin-state $\sigma$ flips, $m_{i}$ changes to $m_{i}-2\sigma$. The energy difference is then:
$\Delta H_{i}=\frac{2J}{\rho_{i}}\left(\sigma m_{i}-1\right)\,.$ (3)
A particle with spin $\sigma$ then flips its state according to the transition rate:
@@
where $\gamma$ is the rate of particle flipping when $\sigma m_{i}=1$. The transition rate is chosen to fulfill the detailed balance without hopping with respect to the Hamiltonian $H_{i}$. In this paper, we choose $J=1$, and $\gamma=1$ without any loss of generality. The parameter $\beta$, denoted as “inverse temperature”, $\beta=T^{-1}$, in passive systems, controls the flip noise strength. Although the system under consideration is athermal, we denote the parameter $T$ as “temperature” from now on.
Moreover, each particle performs a biased diffusion on the lattice depending on the spin state $\sigma$. Particles perform a one-dimensional biased hopping with self-propulsion $\epsilon$ along a direction ${\bf p}$ via the rate Solon and Tailleur (2015):
$W_{\rm hop}(\sigma,{\bf p})=D(1+\sigma\epsilon{\bf p}\cdot{\bf e_{x}})\,.$ (5)
The presence of other particles does not influence hopping rates and hence is independent of particle density. The hopping rate $D=1$ is constant along the upward and downward directions (±$\pm y$-direction). The parameter $\epsilon\in[0,1]$ controls the asymmetry between the purely diffusive limit $\epsilon=0$ and the purely ballistic limit $\epsilon=1$, while $D$ controls the overall hopping rate. On average, a particle drifts with speed $v=2D\epsilon$ in the direction set by the sign of its spin state (where lattice spacing is 1), while the total hopping rate $4D$, remains constant.
@@ III Simulation details
where $f(x)$ is a time-independent scaling function. $R:q(t)$, estimated from the decay of $C(r,t)$ generally show a power-law growth Puri (2009); Bray (2002, 1994):
@@
For scalar order parameters like the density field, the short-distance (large-$k$) behavior of the structure factor scaling function is given by Porod’s law (for domains with smooth boundaries or scattering off sharp domain interfaces) which corresponds to $g(k)\sim k^{-(d+1)}$. Next, we present results for model parameters set by the average particle density $\rho_{0}$, temperature $T=\beta^{-1}$, self-propulsion velocity $\epsilon$, and diffusion constant $D$.
@@ IV.1 Phase diagram and domain morphology
The steady-state behavior of the AIM is summarized in the temperature-density ($T-\rho_{0}$) [Fig. 1(a)] and velocity-density ($\epsilon-\rho_{0}$) [Fig. 1(b)] phase diagrams. The general structure of the phase diagrams consists of a gas phase (G), a liquid phase (L), and a liquid-gas coexistence region (G+L) separated by the gas and liquid binodals. Qualitatively similar results were obtained earlier in Ref. Solon and Tailleur (2015). However, we obtain a critical temperature $T_{c}\simeq 2$ (above which no phase separation occurs regardless of the density, the critical density $\rho_{c}=\infty$) twice as large as in Ref. Solon and Tailleur (2015). This can be understood if one compares the flipping rate of Ref. Solon and Tailleur (2015), $W_{\rm flip}(\sigma\rightarrow-\sigma)=\gamma\exp\left(-\sigma\beta\frac{m_{i}}{\rho_{i}}\right)$ with Eq. (4) for $J=1$. This indicates that the effective $\beta$ considered in this paper is approximately twice as large as the $\beta$ value in Ref. Solon and Tailleur (2015). In the $\epsilon-\rho_{0}$ phase diagram, the two binodals converge at a critical density $(\rho_{c}\simeq 2.9)$ for vanishing self-propulsion ($\epsilon=0$) that signifies a second-order phase transition belonging to the Ising universality class Solon and Tailleur (2013).
@@
To quantify the coarsening dynamics, we performed simulations on a square lattice of size $400^{2}$ with periodic boundary conditions applied on both sides. Following quench at time $t=0$, the system evolves up to $t=10^{5}$ using the MC algorithm described in Sec. III. All numerical data presented here are averaged over at least 300 independent realizations.
Figure 1: (Color online) Phase diagrams of the AIM showing the quench directions at fixed densities. (a) $(T,\rho_{0})$ phase diagram for self-propulsion velocity $\epsilon=1$. (b) $(\epsilon,\rho_{0})$ phase diagram at fixed temperature $T=0.9$. G and L denote the disordered gas and polar liquid regions whereas G+L denotes the phase-coexistence region. The (blue) stars in (a) indicate the initial high-temperature points in the phase diagram from where the system is quenched (indicated by arrows) into the coexistence region ($\rho_{0}=3$) and to the liquid region ($\rho_{0}=10$) [(red) stars at lower temperature]. (Red) stars in (b) mark the quench points in the $(\epsilon,\rho_{0})$ plane. $\rho_{\rm gas}$ (open square, delimits G and G+L) and $\rho_{\rm liquid}$ (open circle, delimits G+L and L) are the gas and liquid binodal, respectively.
@@
A typical simulation of coarsening dynamics starts with a homogeneous initial configuration where particles with $\sigma=\pm 1$ are distributed randomly on each lattice site. This can also be described as an equilibrium configuration at infinite temperature because all configurations are equally likely. Then, after a quench well below the critical temperature ($T_{c}\simeq 2$), the homogeneous initial configuration starts evolving in time, and subsequent dynamics are governed by the formation and growth of $\sigma=1$ and $\sigma=-1$ rich domains. Such a time evolution of the local magnetization field is shown in Fig. 2 for $\epsilon=0$ and $\epsilon=1$. In the latter scenario, time evolution has been shown for a quench in both the coexistence (middle row) and the liquid regions (bottom row). The non-equilibrium steady states (NESS) are characterized by a single band (at lower density) and a polar liquid (at higher density).
@@
The top panel, Fig. 2(a–e), depicts the ordering kinetics of a purely diffusive $(\epsilon=0)$ polar liquid at $\rho_{0}=5$. The domain morphology with increasing time exhibits a close resemblance to the passive Ising model Majumder and Das (2011). In the Ising model, the driving force for domain growth is the curvature of the domain wall, since the system surface energy can only decrease through a reduction in the total net surface area. In the $\epsilon=0$ limit, particles do not form high-density domains due to the diffusive movement of particles. Therefore, density-wise, the system remains homogeneous as we observe a steady growth of small, high-curvature to large, low-curvature domains. This is unsurprising as the $2d$ AIM for $\epsilon=0$ belongs to the same universality class of the passive $2d$ Ising model Solon and Tailleur (2015).
@@
Next, we looked at the evolution of AIM with self-propulsion velocity ($\epsilon=1$) quenching the system inside the spinodal [Fig. 2(f–j)] and homogeneous ordered regions [Fig. 2(k–o)]. The average densities representing these two regions correspond to $\rho_{0}=3$ and 10, respectively. Inside the spinodal region, the growth dynamics is driven by spinodal decomposition and result in the formation of numerous small clusters of negative and positive spins $(t=10^{2})$. The coarsening then stems from the merging of these clusters $(t=10^{3}$ and $t=10^{4})$, until a single macroscopic domain emerges $(t=10^{5})$ in the steady state. A quench deep inside the homogeneous ordered region also results in similar dynamics of cluster formation and coalescence of those clusters into a single large liquid domain (in this paper, for $\epsilon>0$, the word cluster is used interchangeably with domain). With extreme self-propelled particles ($\epsilon=1$), although the high-density clusters are strongly biased along the horizontal directions, they can also grow along the transverse direction due to the constant transverse diffusion $D$. When these clusters merge into a larger cluster, it always tries to minimize the surface energy by decreasing the surface area. Accordingly, the coarsening process leads to a single band [Fig. 2(j)] with domain walls having the lowest curvature (the curvature of a straight line is zero, but a liquid band with a perfectly straight domain boundary can only happen when there is no thermal fluctuation). The focus of this study is therefore to analyze the coarsening dynamics shown in Fig. 2 which we will do next.